OpenStudy (anonymous):

Quick ?.... -4^2 is that a positive 16 or negative 16?

OpenStudy (anonymous):

positive 16

OpenStudy (anonymous):

No, it's negative, actually

OpenStudy (anonymous):

(-4)^2 = 16 -4^2 means -(4^2)

OpenStudy (anonymous):

uber1337h4xx0r is right: (-4)^2=16 but - 4^2=-16

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-4^2 Here's the leading computer program showing it.

OpenStudy (anonymous):

thanks guys.... so i consider it as -4 times -4